OpenStudy (anonymous):

*Just need to know what equation to use* Given: An Ion (q = +2e) enters a magnetic field of 21.5 T at a velocity v perpendicular to the field. If it experiences a force F = 3.58 x 10^-12, determine the magnitude of velocity v of the ion. Thanks

OpenStudy (anonymous):

\[F = q (v \times B) = qvBsin \theta\]

OpenStudy (anonymous):

Thanks! I think I figured it out earlier, if you feel like double checking my answer, that would be cool. I got 5.2 x 10^5 T.

OpenStudy (anonymous):

You're looking for a velocity. I didn't calculate an answer, but it will be in m/s not T.

OpenStudy (anonymous):

yep you are right, I did not mean to add the "T".. I dont even have it on my worksheet haha

OpenStudy (anonymous):

I have another question dealing with the same stuff if you have time.

OpenStudy (anonymous):

An electron is accelerated from rest through a potential difference of 3750 V. It enters a region where B = 4.0 x 10^-3 T perpendicular to its velocity. Calculate the radius of the path it will follow. Would I use the equation: mv^2/r = qvB; where r = mv/qB ?

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

I just checked your answer for the first question and got the same as you.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Yes, your equations look right for the second question.

OpenStudy (anonymous):

Ok, I must be getting confused somewhere. I believe the answer should be 5.2 cm but I cant get to that answer.